OpenStudy (anonymous):

solve the following quadratic equation: 9x^2 - 49 = 0

OpenStudy (nirmalnema):

9x^2 - 49 = 0 (3x)^2 - 7^2 = 0 (use a^2-b^2) (3x-7)(3x+7) x=+_ 7/3

OpenStudy (anonymous):

why is it 7/3?

OpenStudy (nirmalnema):

(3x-7)(3x+7)=0 either (3x-7) =0 or (3x+7)=0

OpenStudy (anonymous):

okay thank you :)